25 Marzo 2014

Observer: F. Cusano, R. Carini, M. Faccini
Telescope Operator: Steve Allanson, Geno Bechetti
Support Astronomer: O. Kuhn (on call)
Instrument: MODS

Summary

Before the sunset the sky was partly cloudy, humidity was at 32% and there was a low wind.
Some problems with modsView and modsAlign in the station obs2, after few check it
seams to be caused by the Ob directory "Antoniucci Bckup".
After the execution of the Antoniucci_bckup OB the sky became clear and the seeing good.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci_bkup V1118Ori dual grating 20m COMPLETED
Gilli_4Q1 J1048 Mask1 red grating 2h COMPLETED
Gilli_4Q2 J1048 Mask2 red grating 2h COMPLETED
Standard Feige34 red grating + dual grating - airmass 1.26 - Default 5" slit
Vanzella C RXJ1532 Mask 4 dual grating 1h  
Vanzella C RXJ1532 Mask 5 dual grating 1h  

26 Marzo 2014

Summary

At the beginnig of the night light SNOW and wind.
At 3:30 no snow but heavy clouds.
At 4:30 completely clear sky, we can't open the dome because of 2-3 inches of snow on the roof. (Mike go up to check the situation)
The TO and John Little agree to stay close and re-check in a couple of hours.

27 marzo 2014

Summary

At the beginnig of the night strong wind 30m/s S-SW stable. Gust around 36m/s.
42% humidity, clear sky!!!
At 6:30 wind is around 22 m/s and the humidity rises to 90%.
The last hours of the night the humidity riches 100 %, dome remains closed.

28 marzo 2014

Summary

Late afternoon the humidity is 102%, outside temperature -2.5 C, dew point temperature -2.2 C, wind around 14 m/s.
Due to the dew point limit, we can open the dome, and the temperatures are : chamber=3.48 C, glass = -0.47 C.
Heavy clouds just on top of the telescope.
the situation go better and we open the dome: temperatures of the glasses are: -0.4 C on the left and -0.8 C on the right, the external is -4 C and the chamber is -3 C.
Problems with the temperature gradient on the mirrors, the focus is not very good.
During the transit we had problems with the blue shutter. It remains stuck opened and we need to stop the ob and operate with the SA every one/two images.
During the transit the seeing value oscillate between 1.4 and 2.9.
We lost about 17 minutes of images during the last part of the transit.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Bellazzini_B HVC204.88+44.86+147 g,r 11m COMPLETED
Bellazzini_A HVC205.28+18.70+150 g,r 11m COMPLETED
Bellazzini_K HVC196.09+24.74+166 g,r 11m to check the last image
Clementini_DDT_2 IZw18 B,r 10m COMPLETED
Nascimbeni_33 gj436 U_spec, F972N20 2.5h COMPLETED , problem with blue shutter in the last hour
Bellazzini_S HVC27.86+38.25+124 g,r 11m COMPLETED
Standard SA101 B, V, g; I, r, r    
Standard SA107 B, V, g; I, r, r   problem with the blue shutter

29 marzo 2014

Summary

We should have used LBC, but the problem with the blue shutter still persists. We use MODS.
Some cirrus at the beginning of the night. Humidity 9%.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Vanzella_D macsj0744 dual grating 1.5h  
Vanzella_E MACSJ1149 dual grating 5 h  
Standard Feige 66 dual grating 6 m  
Vanzella_C RXJ1532 dual grating 0.5 h  

-- %USERSIG{RobertaCarini - 2014-03-29}%
Topic revision: r2 - 29 Mar 2014, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback